# Quantum cryptography beyond key distribution: theory and experiment

### v1

@@ IV.6 One-time programs
One-time programs (OTPs) were introduced as a computational paradigm for security applications, ranging from software protection and program obfuscation to the temporary transfer of cryptographic abilities Goldwasser et al. (2008). Ideally, an OTP acts as a black box function which can only be evaluated on a single input before self-destructing. This implies that a user cannot access the inner workings of the program besides a single input-output relation.
@@ V Security based on other physical assumptions
This section discusses how additional physical assumptions may be combined with conjugate coding to provide i.t.-secure implementations of mistrustful primitives, thus circumventing this no-go result. In a nutshell, the first two approaches enforce that a dishonest party measures their quantum states at a given point in time, either by exploiting no-signalling constraints or by making assumptions on the size/noise of the adversary’s quantum storage device. The third approach relies on physical unclonable hardware assumptions.
@@ V.1 Special relativity and trusted agents
This protocol was later refined and experimentally demonstrated in Lunghi et al. (2013); Liu et al. (2014). From a practical standpoint, it is crucial to note that the commitment time is a function of distance $d$, thus imposing large distances between a party’s trusted agents. Locating the agents in Singapore and Geneva, for example, provides a commitment time of $15$ ms Lunghi et al. (2013). On the security side, one must consider that monitoring the spatial and temporal coordinates of verifiers requires a trusted Global Positioning System (GPS), which might open the door to spoofing attacks and time-stamp manipulation Tippenhauer et al. (2011). This may however be mitigated if secure position verification schemes are used (Section III.3). Furthermore, the composability of such relativistic protocols must be carefully assessed Vilasini et al. (2019).
@@ V.3 Physical unclonable functions
The first quantum approach was introduced in Škorić (2010), where the concept of quantum readout PUF (QR-PUF) was defined. The idea is that a quantum challenger prepares a superposition of challenges and receives back the corresponding superposition of responses. Any party intercepting the quantum states would be essentially detected, thus ensuring that a challenger is indeed probing a specific QR-PUF, even when it is under adversarial control. The security here relies on the principles of quantum theory. A variant involving coherent states with low mean photon number was later implemented Goorden et al. (2014), while another protocol, based on continuous variable with homodyne detection, was demonstrated in Nikolopoulos and Diamanti (2017). Furthermore, a theoretical framework for QR-PUFs was developed in Gianfelici et al. (2020). While offering interesting prospects, these protocols all presented two main disadvantages. First, their security was proven against specific types of attacks (e.g. Nikolopoulos (2018)), but not in full generality. Second, the verifier requires some information on the quantum channel implemented by the PUF, typically in order to perform the quantum readout. This differs from the initial classical notion of PUFs.
@@ VI Computational security against quantum attacks
Currently, there are two main approaches: the first approach, which is not covered in this review, uses classical resources to design primitives that are secure against quantum attacks. Labelled post-quantum cryptography, this field is evolving very rapidly, both in terms of standardization (through the NIST competition) and cryptanalysis, in which security conjectures are broken and new schemes appear. The second approach, which is the main focus of this section, uses quantum resources to protect against quantum attacks, unlocking new protocols such as public-key quantum money, one-shot signatures, public-key encryption and zero-knowledge proofs. For completeness, we provide a very brief discussion of cryptanalysis in Section VI.5.
@@ VI.2 Tokenized and one-shot signatures
Despite these exciting applications, two main challenges stand in the way of one-shot signatures being practically realized. The first one, that is shared with public-key quantum money, is the hardware requirements, as it requires a large scale (to have sufficient qubits), fault-tolerant (to have sufficient depth), quantum computation in order to resist brute-force attacks. The second issue relates to the practicality, and even the existence, of suitable hash functions satisfying all the theoretical requirements to guarantee security, which remains an open question.
@@ VI.3 Public-key cryptography
The notion of public-key cryptography was introduced in Diffie and Hellman (1976) and first implemented by RSA Rivest et al. (1978). A public-key encryption (PKE) scheme comprises three efficient algorithms: $\operatorname{Gen,Enc,Dec}$. The first one, $\operatorname{Gen}$, is used to generate a pair of secret and public keys $sk,pk$. The public key is published for everyone to use while the secret key is kept secret by the creator of the system. The second algorithm, $\operatorname{Enc}$, maps the public key $pk$ and message $m$ to a ciphertext $c$. Finally, the third algoritm $\operatorname{Dec}$ maps the secret key $sk$ and ciphertext $c$ back to message $m$.
The great advantage of PKE is that no secret exchange between sender and receiver is necessary to send encrypted messages. It suffices for the sender to get hold of the public key associated with the receiver. In the classical setting, PKE can be obtained from trapdoor one-way functions Diffie and Hellman (1976). In the classical setting, it is simple to understand why i.t.-secure PKE is impossible: if $pk$ determines $sk$ completely, then any receiver can compute $sk$ and decrypt every possible ciphertext $c$.
Interestingly, it was shown that i.t.-secure quantum public-key encryption (qPKE) is also impossible Barooti et al. (2023). Due to the public key being a quantum state, this impossibility is much less obvious than for the classical case. In fact, some of the existing constructions of qPKE had been conjectured to be unconditionally secure Gottesmann (2005), a conjecture that is invalidated by the work from Barooti et al. (2023).
@@ VI.5 Cryptanalysis
Due to Shor’s algorithm Shor (1997), we know that the most commonly used cryptosystems will break when a sufficiently large fault-tolerant quantum computer is available. Cryptanalysis, however, deals with specific numbers that arise from the parameters of each cryptosystem, and in this case, we need to determine what resources would be needed to actually break standard cryptosystems such as RSA Rivest et al. (1978). This of course depends on many parameters. Indicatively: the most optimised quantum factoring algorithm, the noise level and speed of gates of the quantum hardware and the quantum error-correcting code used. A very thorough analysis was provided in Gidney and Ekerå (2021), where the authors show that RSA 2048 can be broken in eight hours using a superconducting quantum computer with 20 million qubits of the same quality (fidelity, coherence, gate-speed) as those from concurrent papers.
@@ VII.1 Secret sharing
The motivation of the first quantum scheme was to render eavesdropping attempts on both strings sent by Alice detectable Hillery et al. (1999). In the honest scenario, each of the three parties holds one subsystem of a $3$-partite GHZ state and performs a measurement in a basis uniformly chosen from $\{X,Y\}$. This is repeated for many GHZ states. After all three parties publicly declare their measurement bases (but not the outcomes), Bob and Charlie may combine their outcome strings to infer Alice’s string in half of the cases. Similarly to QKD, the presence of eavesdropping (whether from an external party or from Bob/Charlie intercepting Charlie/Bob’s subsystem) may be tested by public comparison of a subset of their measurements.
@@
Given the low count rates achievable for $3$- and $4$-photon entanglement at the time Pan et al. (2001); Bouwmeester et al. (1999), experimental quantum secret sharing was first demonstrated using pseudo-GHZ states, in which $3$-partite correlations were extracted from two down-converted photons and one pump photon instead of three down-converted photons Tittel et al. (2001). A full scheme was later demonstrated using $4$-qubit polarization GHZ states, in which one of the photons is used to herald a successful $3$-partite GHZ state with visibility up to $83\%$ Chen et al. (2005), followed by a $4$-partite scheme Gaertner et al. (2007). As an alternative to multipartite entangled states, a scalable "single photon" alternative was demonstrated between $6$ parties, in which each party sequentially imprints a phase on a single-photon that they forward to the next party Schmid et al. (2005). This was later extended to a travelling single qudit experiment Smania et al. (2016). Finally, a protocol using phase encoding of weak coherent states was demonstrated Wang et al. (2024), along with continuous-variable implementations Zhou et al. (2018); Liu et al. (2023a); Qin et al. (2024) and secret sharing involving quantum secrets Lu et al. (2016); Bell et al. (2014).
@@ VII.3 Byzantine agreement
With the goal of photonic demonstration in mind, cheat-sensitive schemes using more practical resources were designed, replacing the tripartite entangled qutrit state from Eq. (15) with $4$-partite entangled qubit states Cabello (2003), multipartite entangled states in the continuous variable setting Neigovzen et al. (2008); Feng et al. (2019), $3$-partite GHZ states Qu et al. (2023) and trusted Bell pair sources Andronikos and Sirokofskich (2023).
After the first photonic implementation was provided using a $4$-photon polarization-entangled state Gaertner et al. (2008), a proof-of-concept version was implemented using a single (i.e. non-entangled) qutrit state, on which each party imprints a phase before forwarding it to the next party Smania et al. (2016). The qutrit’s three levels are encoded onto spatial modes in a $3$-arm Mach-Zehnder-type interferometer. After the last party measures in the Fourier basis, all parties can check for consistency by declaring their respective applied phases. This setup has the interesting benefit that the number of single-photon detectors does not scale with the number of parties $n$. Very recently, an i.t.-secure implementation beating the $n>3k$ classical limit was performed using quantum digital signatures based on one-time universal hashing Li et al. (2023), an integrated polarization-entangled photon pair source and optical fiber spools of a few kilometers Jing et al. (2024).
@@ VIII.1 Private channels
An alternative encryption scheme uses classical keys and produces a quantum ciphertext Ambainis et al. (2000); Boykin and Roychowdhury (2003). If |$\ket{\psi}$ is a qubit to be encrypted, then the encryption scheme uses a pair of random bits $(a,b)$ as the classical key and |$\ket{\psi_{ciph}}=X^{a}Z^{b}\ket{\psi}$ as the quantum ciphertext. For any party who is completely ignorant about the key, the mixed state associated with |$\ket{\psi_{ciph}}$ is indeed the completely mixed state. Since the latter is independent of the input |$\ket{\psi}$, this scheme is also perfectly secret in the sense of Shannon, and known as the quantum Vernam cipher.
@@ VIII.3 Anonymous communication
While some primitives ensure the confidentiality of messages, anonymity ensures the confidentiality of identity. In scenarios such as online auctions, one party may wish to broadcast a message to a set of network users while keeping their own identity secret Stajano and Anderson (2000). Using classical resources, i.t.-secure anonymous broadcasting is possible assuming pairwise authentic private channels and reliable broadcast channels, as in the Dining Cryptographers Problem Chaum (1988), or using trusted servers. Depending on the exact application, some protocols may only require sender or receiver anonymity, while others may require both (full anonymity).
@@ IX.1 Blind computation
The first scenario involves one server with a (universal) quantum computer and one client with limited computational capabilities. The client does not trust the server but wishes to delegate their computation to them. In most cases, we assume that the client can only prepare and send single qubits to the server. This setting was first termed blind quantum computing (BQC) in Arrighi and Salvail (2006), where a first, limited, example of delegation of some quantum computation was given. The first protocol that allowed to delegate universal quantum computation with minimal resources was introduced in Broadbent et al. (2009).
@@
where the angle $\theta$ for each position is chosen uniformly at random (and independently) by the client and kept secret from the server. The protocol uses measurement-based quantum computing (MBQC), which is a universal model of quantum computation where the computation is performed by consecutive single-qubit measurements, where the outcome of the previous measurement defines the next measurement setting, on a generic large entangled “resource” state, independent of the computation Briegel et al. (2009); Wei (2021)). The single-qubit measurement angles $\phi$ determine the unitary an MBQC “pattern” implements.
@@
Figure 7: Universal blind quantum computing. First proposed in Broadbent et al. (2009), this figure was taken from Gheorghiu et al. (2019), and reproduced with permission from the authors. The client prepares a stream of random qubits of the form |$\ket{+_{\theta}}$, sends them to the server, and instructs them to place them at a grid to prepare a rotated graph state. The client instructs the server to measure the qubits in the basis corresponding to a $\delta$ angle that cancels the pre-rotation and performs the desired computation in a blind way.

### v2

@@ IV.6 One-time programs
One-time programs (OTPs) were introduced as a computational paradigm for security applications, ranging from software protection and program obfuscation to the temporary transfer of cryptographic abilities Goldwasser et al. (2008). Ideally, an OTP acts as a black box function which can only be evaluated on a single input before self-destructing. This implies that a user cannot access the inner workings of the program besides a single input-output relation. More formally, the OTP cryptographic primitive for a function $f$ ensures that an efficient receiver can evaluate $f(x)$ at run time, but cannot learn anything about $f(y)$ for $y\neq x$ beyond what they can learn from $f(x)$.
@@ V Security based on other physical assumptions
This section discusses how the additional physical assumptions summarized in Table 1 may be combined with conjugate coding to provide i.t.-secure implementations of mistrustful primitives, thus circumventing this no-go result. In a nutshell, the first two approaches enforce that a dishonest party measures their quantum states at a given point in time, either by exploiting no-signalling constraints or by making assumptions on the size/noise of the adversary’s quantum storage device. The third approach relies on physical unclonable hardware assumptions.
@@ V.1 Special relativity and trusted agents
This protocol was later refined and experimentally demonstrated in Lunghi et al. (2013); Liu et al. (2014). From a practical standpoint, it is crucial to note that the commitment time is a function of distance $d$, thus imposing large distances between a party’s trusted agents. Locating the agents in Singapore and Geneva, for example, provides a commitment time of $15$ ms Lunghi et al. (2013). By designing multi-round classical protocols, arbitrary long commitment times secure against classical attacks are in principle possible Chakraborty et al. (2015); Lunghi et al. (2015), and were experimentally demonstrated up to $24$ hours Verbanis et al. (2016). Generally, one must consider that monitoring the spatial and temporal coordinates of verifiers requires a trusted Global Positioning System (GPS), which might open the door to spoofing attacks and time-stamp manipulation Tippenhauer et al. (2011). This may however be mitigated if secure position verification schemes are used (Section III.3). Furthermore, the composability of such relativistic protocols must be carefully assessed Vilasini et al. (2019).
@@ V.2 Bounded and noisy storage
More recently, i.t.-secure constructions under the bounded storage model were proposed for asymmetric key encryption and signature schemes Barhoush and Salvail (2025).
@@
As an alternative or addition to quantum storage assumptions, we note the existence of BC schemes in which the emission time of each quantum state is randomized by the sender, which yields secure protocols under the assumption that a dishonest receiver cannot perform quantum non-demolition measurements Danan and Vaidman (2011); Loura et al. (2014, 2016).
@@ V.3 Physical unclonable functions
The first quantum approach was introduced in Škorić (2010), where the concept of quantum readout PUF (QR-PUF) was defined. The idea is that a quantum challenger prepares a superposition of challenges and receives back the corresponding superposition of responses. Any party intercepting the quantum states would be essentially detected, thus ensuring that a challenger is indeed probing a specific QR-PUF, even when it is under adversarial control. The security here relies on the principles of quantum theory. A variant involving coherent states with low mean photon number was later implemented Goorden et al. (2014), while another protocol, based on continuous variable with homodyne detection, was demonstrated in Nikolopoulos and Diamanti (2017). Furthermore, a theoretical framework for QR-PUFs was developed in Gianfelici et al. (2020). While offering interesting prospects, these protocols all present two main disadvantages. First, their security is proven against specific types of attacks (e.g. Nikolopoulos (2018); Fladung et al. (2019)), but not in full generality. Second, the verifier requires some information on the quantum channel implemented by the PUF, typically in order to perform the quantum readout. This differs from the initial classical notion of PUFs.
@@ VI Computational security against quantum attacks
Currently, there are two main approaches: the first approach, which is not covered in this review, uses classical resources to design primitives that are secure against quantum attacks. Labelled post-quantum cryptography, this field is evolving very rapidly, both in terms of standardization (through the NIST competition) and cryptanalysis, in which security conjectures are broken and new schemes appear. The second approach, which is the main focus of this section, uses quantum resources to protect against quantum attacks, unlocking new protocols summarized in Table 1 such as public-key quantum money, one-shot signatures, public-key encryption and zero-knowledge proofs. For completeness, we provide a very brief discussion of cryptanalysis in Section VI.5.
@@ VI.1 Public-key quantum money
Recently, a first step towards noise-tolerant public-key quantum money was proposed, along with its security challenges Yuen (2025). Essentially, the subspace state from Eq. (12) may either be treated as a codeword in a Calderbank-Shor-Steane code, or encoded and decoded using an additional layer of quantum error correction.
@@ VI.2 Tokenized and one-shot signatures
Despite these exciting applications, two main challenges stand in the way of one-shot signatures being practically realized. The first one, that is shared with public-key quantum money, is the hardware requirements, as it requires a large scale (to have sufficient qubits), fault-tolerant (to have sufficient depth), quantum computation in order to resist brute-force attacks. The second issue relates to practicality. Recently, a concrete scheme was proposed to realize the functionality, relying solely on quantum secure indistinguishability obfuscation (iO) and the hardness of Learning with Errors, thus resolving the theoretical question of whether OSS are possible Shmueli and Zhandry (2025). However, the said construction would lead to very high overheads if one wanted to implement it, meaning that further research on reducing the complexity and cost is essential before being realized in practice.
@@ VI.3 Public-key cryptography
The notion of public-key cryptography was introduced in Diffie and Hellman (1976) and first implemented by RSA Rivest et al. (1978). Besides the latter, public-key encryption algorithms include ElGamal Elgamal (1985) based on discrete logarithm or elliptic curves Koblitz (1987), Cramer-Shoup Cramer and Shoup (1998) and lattice-based key encapsulation recently standardized by NIST NIS (2024).
A public-key encryption (PKE) scheme comprises three efficient algorithms: $\operatorname{Gen,Enc,Dec}$. The first one, $\operatorname{Gen}$, is used to generate a pair of secret and public keys $sk,pk$. The public key is published for everyone to use while the secret key is kept secret by the creator of the system. The second algorithm, $\operatorname{Enc}$, maps the public key $pk$ and message $m$ to a ciphertext $c$. Finally, the third algoritm $\operatorname{Dec}$ maps the secret key $sk$ and ciphertext $c$ back to message $m$.
The great advantage of PKE is that no secret exchange between sender and receiver is necessary to send encrypted messages. It suffices for the sender to get hold of the public key associated with the receiver. In the classical setting, PKE can be obtained from trapdoor one-way functions Diffie and Hellman (1976), and it is simple to understand why i.t.-secure PKE is impossible: if $pk$ determines $sk$ completely, then any receiver can compute $sk$ and decrypt every possible ciphertext $c$.
Interestingly, it was shown that i.t.-secure quantum public-key encryption (qPKE) is also impossible Nikolopoulos (2008); Nikolopoulos and Ioannou (2009); Seyfarth et al. (2012); Barooti et al. (2023). Due to the public key being a quantum state, this impossibility is much less obvious than for the classical case. In fact, some of the existing constructions of qPKE had been conjectured to be unconditionally secure Gottesmann (2005), a conjecture that is invalidated by the work from Barooti et al. (2023).
@@ VI.4 Zero-knowledge proofs
Two experimental demonstrations of zero-knowledge proofs have been proposed for the problem of graph $3$-coloring. The first aims for i.t.-security by relying on multiple provers that cannot communicate with each other Alikhani et al. (2021), harnessing the relativistic assumptions of Section V.1. The second uses device-independent quantum randomness and a hash-based signature scheme to construct a non-interactive zero-knowledge proof without resorting to the random oracle model Li et al. (2023c).
@@ VI.5 Cryptanalysis
Due to Shor’s algorithm Shor (1997), we know that the most commonly used cryptosystems will break when a sufficiently large fault-tolerant quantum computer is available. Cryptanalysis, however, deals with specific numbers that arise from the parameters of each cryptosystem, and in this case, we need to determine what resources would be needed to actually break standard cryptosystems such as RSA Rivest et al. (1978). This of course depends on many parameters. Indicatively: the most optimised quantum factoring algorithm, the noise level and speed of gates of the quantum hardware and the quantum error-correcting code used. A very thorough analysis was provided in Gidney and Ekerå (2021), where the authors show that RSA 2048 can be broken in eight hours using a superconducting quantum computer with 20 million qubits of the same quality (fidelity, coherence, gate-speed) as those from concurrent papers. However, such estimates can significantly change, purely from theoretical developments in quantum error correction and quantum algorithms optimization. In a very recent follow-up work, and making the same assumptions on the quality of qubits, Gidney (2025) used three novel methods to reduce the required physical qubits to break RSA 2048 to under one million – giving a twenty-fold improvement over prior state-of-the-art.
@@ VII.1 Secret sharing
The motivation behind the first quantum schemes was to render eavesdropping attempts on both strings sent by Alice detectable Hillery et al. (1999); Żukowski et al. (1998). In the honest scenario, each of the three parties holds one subsystem of a $3$-partite GHZ state and performs a measurement in a basis uniformly chosen from $\{X,Y\}$. This is repeated for many GHZ states. After all three parties publicly declare their measurement bases (but not the outcomes), Bob and Charlie may combine their outcome strings to infer Alice’s string in half of the cases. Similarly to QKD, the presence of eavesdropping (whether from an external party or from Bob/Charlie intercepting Charlie/Bob’s subsystem) may be tested by public comparison of a subset of their measurements.
@@
Given the low count rates achievable for $3$- and $4$-photon entanglement at the time Pan et al. (2001); Bouwmeester et al. (1999), experimental quantum secret sharing was first demonstrated using pseudo-GHZ states, in which $3$-partite correlations were extracted from two down-converted photons and one pump photon instead of three down-converted photons Tittel et al. (2001). A full scheme was later demonstrated using $4$-qubit polarization GHZ states, in which one of the photons is used to herald a successful $3$-partite GHZ state with visibility up to $83\%$ Chen et al. (2005), followed by a $4$-partite scheme Gaertner et al. (2007). As an alternative to multipartite entangled states, a scalable "single photon" alternative was demonstrated between $6$ parties, in which each party sequentially imprints a phase on a single-photon that they forward to the next party Schmid et al. (2005). This was later extended to a travelling single qudit experiment Smania et al. (2016).
Alongside phase-encoded proposals for quantum secret sharing in the measurement device-independent setting Fu et al. (2015); Li et al. (2023b), a weak coherent state scheme was demonstrated Wang et al. (2024). Finally, it is worth noting the implementation of continuous-variable schemes Zhou et al. (2018); Liu et al. (2023a); Qin et al. (2024) and secret sharing protocols involving quantum secrets Lu et al. (2016); Bell et al. (2014). For a more detailed discussion of quantum primitives aiming to protect quantum information, we refer the Reader to Section VIII.
@@ VII.3 Byzantine agreement
With the goal of photonic demonstration in mind, cheat-sensitive schemes using more practical resources were designed, replacing the tripartite entangled qutrit state from Eq. (15) with $4$-partite entangled qubit states Cabello (2003); Guba et al. (2024), multipartite entangled states in the continuous variable setting Neigovzen et al. (2008); Feng et al. (2019), $3$-partite GHZ states Qu et al. (2023) and trusted Bell pair sources Andronikos and Sirokofskich (2023).
After the first photonic implementation was provided using a $4$-photon polarization-entangled state Gaertner et al. (2008), a proof-of-concept version was implemented using a single (i.e. non-entangled) qutrit state, on which each party imprints a phase before forwarding it to the next party Smania et al. (2016). The qutrit’s three levels are encoded onto spatial modes in a $3$-arm Mach-Zehnder-type interferometer. After the last party measures in the Fourier basis, all parties can check for consistency by declaring their respective applied phases. This setup has the interesting benefit that the number of single-photon detectors does not scale with the number of parties $n$. Recently, an i.t.-secure Byzantine agreement framework breaking the $n>3k$ classical limit was developed by using a decentralized three-party relationship of quantum digital signatures (Section IV.1) to eliminate channel independence, along with a recursive structure to extend it to multipartite scenarios Weng et al. (2023). By utilizing an integrated polarization-entangled photon pair source and optical fiber spools of several kilometers, this framework was experimentally demonstrated Jing et al. (2024), employing the one-time universal hashing signature scheme from Li et al. (2023a).
@@ VIII.1 Private channels
An alternative encryption scheme uses classical keys and produces a quantum ciphertext Ambainis et al. (2000); Boykin and Roychowdhury (2003). If |$\ket{\psi}$ is a qubit to be encrypted, then the encryption scheme uses a pair of random bits $(a,b)$ as the classical key and |$\ket{\psi_{ciph}}=X^{a}Z^{b}\ket{\psi}$ as the quantum ciphertext. For any party who is completely ignorant about the key, the security guarantee is that the mixed state associated with |$\ket{\psi_{ciph}}$ is indeed the completely mixed state, i.e. appears random to that party. Since the latter is independent of the input |$\ket{\psi}$, this scheme is also perfectly secret in the sense of Shannon, and known as the quantum Vernam cipher.
@@ VIII.2 Authentication
Informally, we consider a MAC to be secure if existential unforgeability against a chosen message attack holds, where:
  1. (i)
    existentially unforgeable means that the attacker is not able to generate a valid MAC tag on any message, without being in possession of the key $k$.
  2. (ii)
    chosen message attack means that the attacker is able to obtain MAC tags on any other message before performing the attack.
@@ VIII.3 Anonymous communication
While some primitives ensure the confidentiality of messages, anonymity ensures the confidentiality of identity. In scenarios such as online auctions, one party may wish to broadcast a message to a set of network users while keeping their own identity secret Stajano and Anderson (2000). Using classical resources, i.t.-secure anonymous broadcasting is possible assuming pairwise authentic private channels and reliable broadcast channels, as in the Dining Cryptographers Problem Chaum (1988), or using trusted servers. Depending on the exact application, some protocols may only require sender or receiver anonymity, while others may require both (full anonymity). Informally, anonymity for a sender (resp. receiver) is guaranteed if the communication taking place throughout the protocol does not change the a priori uncertainty about the identity of the sender (resp. receiver).
@@ IX.1 Blind computation
The first scenario involves one server with a (universal) quantum computer and one client with limited computational capabilities. The client wishes to delegate their computation in such a way that their input, output and computation remain perfectly hidden from the server. In other words, we have that
  1. (i)
    the distribution of classical information obtained by the server during the client’s instructions and from the output of the computation
  2. (ii)
    the quantum state of the systems received by the server during the protocol
must appear independent of the client’s input.
In most cases, we assume that the client can only prepare and send single qubits to the server. This setting was first termed blind quantum computing (BQC) in Arrighi and Salvail (2006), where a first, limited, example of delegation of some quantum computation was given. The first protocol that allowed to delegate universal quantum computation with minimal resources was introduced in Broadbent et al. (2009).
@@
where the angle $\theta$ for each position is chosen uniformly at random (and independently) by the client and kept secret from the server. The protocol uses measurement-based quantum computing (MBQC), which is a universal model of quantum computation in which the computation is performed by consecutive single-qubit measurements and the outcome of the previous measurement defines the next measurement setting, on a generic large entangled “resource” state that is independent of the computation Briegel et al. (2009); Wei (2021)). The single-qubit measurement angles $\phi$ determine the unitary implemented by an MBQC “pattern”.
@@
Figure 7: Universal blind quantum computing. First proposed in Broadbent et al. (2009), this figure was taken from Gheorghiu et al. (2019), and reproduced with permission from the authors. For each position $(i,j)$ in the rectangular cluster state, the client prepares a random qubit of the form |$\ket{+_{\theta}}_{ij}$ as defined in Eq. (17), sends it to the server, and requires them to prepare the rotated graph state. The client then instructs the server to perform a measurement $M(\delta_{ij})$ in a basis with angle $\delta_{ij}$, which cancels the pre-rotation $\theta_{ij}$ and performs the desired computation dictated by $\phi_{ij}$ in a blind way. Note that the extra term $r_{ij}\pi$ essentially one-time pads the measurement outcome $b_{ij}$.